Use the law of reflection to determine the minimum height and position of a plane mirror that just allows a 5'6" woman standing on the floor in front of the mirror to Sense both her head and Feet in the mirror
1
Expert's answer
2021-08-23T12:38:32-0400

"\text{Let AB segment denoting the position of a woman.}"

"AB = 5'6''"

"A(0,5'6'') \text{ -point head position}"

"B(0,0)\text{- point leg position}"

"E(0,E_y)\text{-point eye position}"

"CD\text{ segment indicating the position of the mirror}"

"\text {Let's draw the bisector }CT \vartriangle ACE"

"СT \text{ is also the height of the triangle, therefore}"

"C_y = A_y-\frac{AE}{2}"

"\text {Let's draw the bisector }DL \vartriangle BDE"

"DL \text{ is also the height of the triangle, therefore}"

"D_y = \frac{BE}{2}"

"\text{The maximum possible eye point is point } A"

"\text{then}"

"C_y = A_y=5'6''"

"D_y = \frac{AB}{2}= 2''9'"

"CD = 2''9'"


"\text{Answer: mirror length -}2''9'"

"\text{mirror height above floor level - }2''9'"